NEXT
Make New Entry,
Make Followup Entry
User name jroche
Log entry time 08:15:27 on October 12, 2010
Entry number 334494
keyword=Owl shift summary
October 11, 2010, Owl shift summary
Crew:
J. Roche
F. Itard
V. Bellini
Timeline:
-
00:00: start night with run 7205 in progress
: LH2 target, no sieve, pnom=3.194 GeV (Elastic -4%), theta=27.5 deg
: beam position: A(x,y):-0.3, 0.3, B(x,y):0,0, 0.4. J. Huan reports
that trying for a better cntering get the ions chamber to trip. We leave
the beam position as is.
: MCC is trouble shooting the FFB that cannot be locked on both Hall A
and C atthe same time. FFB is off.
: data acquisition rate= 71 Hz at 10 uA and with the following prescales:
ps1=0,ps2=0,ps3=1,ps4=1,ps5=0,ps6=0,ps7=0,ps8=1000,ps9=0
- M. Friend is in the backroom looking at the Compton (
here).
- 1:50: Done with Elastic optics without sieve slit. We are taking a
short run with ps1 trigger only to check timing spectra with the S2 as a
trigger only: run 7207.
- 2:00: changed target to Carbon foils + change magnet momentum setting
- 3:00: request 50 uA MCC will only give us 25 uA, Abort Carbon foils
program.
- 3:22: put LH2 15 cm in, put sieve in position.
- 3:33: request 10 uA on LH2.Rates with sieve=38 Hz
- 4:53: instabilities in the mchine, MCC tasks the beam away.
- 5:23: beam is back. MCC had to calibrate the Hall C bcms as Hall C
was finally taking significant beam (20uA)
- 7:35: MCC takes beam away to do a QE scan
- Vincenzo notices that the running conditions have changed without
apparent reasons. The trigger rates went up and we lost the pattern due
to the sieve slit.
Run list:
ps1=0,ps2=0,ps3=1,ps4=1,ps5=0,ps6=0,ps7=0,ps8=1000,ps9=0
- 7205: LH2 target 15cm, no sieve, pnom=3.194 GeV (Elastic -4%),
theta=27.5 deg, sieve out, 541 k events
- 7206: junk
ps1=25,ps2=0,ps3=0,ps4=0,ps5=0,ps6=0,ps7=0,ps8=0,ps9=0
- 7207: LH2 target 15cm, no sieve, pnom=3.194 GeV (Elastic -4%),
theta=27.5 deg, sieve out, 320 k events
ps1=0,ps2=0,ps3=1,ps4=1,ps5=0,ps6=0,ps7=0,ps8=1000,ps9=0
- 7208: Carbon optics, pnom=3.327 GeV (Elastic 0%), theta=27.5 deg,
sieve out, 25 uA. Very low rate: 29 Hz.
- 7209: LH2, pnom=3.327 GeV (Elastic 0%), theta=27.5 deg. Sieve in.
- 7210: LH2, pnom=3.327 GeV (Elastic 0%), theta=27.5 deg. Sieve in.
- 7211: LH2, pnom=3.327 GeV (Elastic 0%), theta=27.5 deg. Sieve in.